OpenStudy (jamanorthi):

how do i simplify 3x(5x-2) plz help!!!

OpenStudy (bibby):

Distribute the 3x to both terms like so: 3x(5x-2) = (3x*5x) - (3x*2)

OpenStudy (jamanorthi):

so it 15x-6 rigth?

OpenStudy (bibby):

Not exactly. x*x = x^2 not x

OpenStudy (jamanorthi):

so what it gives?

OpenStudy (bibby):

I don't know. Learn to multiply. 3x(5x-2) = (3x*5x) - (3x*2)

OpenStudy (jamanorthi):

yeah its 15x-6

OpenStudy (bibby):

if you say so. I'll take your word for it as I don't know math

OpenStudy (jamanorthi):

okay bye